OpenStudy (anonymous):

An inlet tap A inlets water at 2 litre/min and another tap B inlets water at 3 litre/min. There is an outlet pipe at the bottom, which can empty the tank in 10 hours. If all the three taps were kept on, the tank had been filled in 8 hours. What is the capacity of the tank?

OpenStudy (anonymous):

volume=rate*time. let r be the rate of pipe C v=r*10 (There is an outlet pipe at the bottom, which can empty the tank in 10 hours) v=(2+3-r)*8 (If all the three taps were kept on, the tank had been filled in 8 hours) solve for v

OpenStudy (anonymous):

no we have another variable r

OpenStudy (anonymous):

10r=40-8r==>r=20/9

OpenStudy (anonymous):

1/2+1/3+1/x=1/8, x=17/24,

OpenStudy (whpalmer4):

Rate for A is 2 liters/min Rate for B is 3 liters/min Total inflow is 5 liters/min =300 liters/hr Outflow is v/10 liters/hr where v is volume of tank (Inflow-outflow)*8 = v

OpenStudy (anonymous):

24000=18v,v=1333.33liters

OpenStudy (whpalmer4):

Yep. Check answer: outflow is 133.3 l/hr so net inflow is 300-133.3=166.7 l/hr 166.7*8=1333.6

OpenStudy (whpalmer4):

If we left answer in fractional form check would be exactly correct.